having problem solving to algebra problems
1. 18 + (5/6) r=0


2. 3x + (2/3) = 5/6

Answer by nerdybill(7384)

1. 18 + (5/6) r=0
Multiply both sides by 6:
108 + 5r = 0
5r = -108
r = -108/5
r = -21.6
2. 3x + (2/3) = 5/6
Multiply both sides by 6
18x + 4 = 5
18x = 1
x = 1/18
x = 0.056